# Modular Compilation for Quantum Chiplet Architectures

### v3

@@ 5. Experimental Results
Figure 13 shows the recurring compilation time. Specifically, this refers to the elaboration time of the SEQC compilers and the full compilation time for Qiskit and PA-Qiskit. Overall, SEQC compilers have a large speedup over Qiskit-based compilers, owing to the use of parallelism and modular compilation strategies. Compared to Qiskit, SEQC-STA achieves 3.27×3.27$\times$3.27 × speedup on average (up to 6.74×6.74$\times$6.74 ×), and compared to PA-Qiskit 4.69×4.69$\times$4.69 × average speedup (6.61×6.61$\times$6.61 × max).
@@ 6. Discussion
These results demonstrate a flaw with gate count and circuit depth as heuristics: their lack of noise-awareness. Circuit depth and gate count only consider the number of gates (in the critical path and in total) irrespective of each one’s unique underlying properties. It is expected that two-qubit gates will be significantly more noisy than single-qubit gates, and inter-chiplet gates should be significantly more noisy than intra-chiplet gates (smith_2022_chiplets). Failure to account for duration and error differences between gates or types of gates in quantum compilation heuristics will inevitably result in sub-optimal solutions when the discrepancies are extreme, e.g., in chiplet architectures. Conversely, the hierarchical structure of the SEQC framework with the stratification and elaboration stages imparts some degree of implicit noise-awareness, where minimizing the number of inter-chiplet gates takes precedence over intra-chiplet/total gates. As a result, SEQC can better optimize for the practically significant metrics of circuit fidelity and circuit execution time.
Overall, we observe that SEQC yields circuits with more desirable characteristics when executing on modular quantum architectures, while at the same time it achieves faster compilation times. Also, as the number of qubits increases, the performance improvement of SEQC widens compared to the baseline, especially for the most important metrics of fidelity, compiled circuit execution time, and number of inter-chiplet gates. These trends hold promise in allowing for faster and better compilation for future large-scale modular quantum devices than we can achieve with today’s frameworks.
@@ 7. Related Work
MECH (mech) sacrifices some ancillary qubit resources to build an ”adjustable multi-entry communication highway” through software. By entangling ancillary qubits and distributing them throughout a chiplet, MECH virtualizes those qubits into fast network-on-chip routing channels. Lin et al. (let_every_qubit) present a compilation technique that utilizes calibration data to determine potentially nonstandard 2Q gates that can perform standard operations along bases more compatible with their respective physical states and improve fidelity. Bandic et al.(bandic2024profiling) utilize the relationship between quantum circuit structure and mapping efficiency to optimize the architecture of monolithic and modular quantum processors, on the path to application-specific hardware optimization. Escofet et al. (escofet2024revisiting) develop a characterization technique for arbitrary quantum circuits which yields theoretical upper and lower optimization bounds, enabling the optimality assessment of qubit mapping solutions. All these techniques are orthogonal to SEQC, and could work synergistically with it to achieve the combined benefits.
@@
CutQC (cutqc) enables smaller-scale quantum computers to execute quantum circuits that are otherwise too large to fit on the device. It partitions large quantum circuits into smaller subcircuits, executes them one at a time, and then merge the individual results at the cost of classical post-processing. Piveteau and Sutter (circuit_knitting) similarly allow subcircuits to execute independently, and use quasiprobability simulation to simulate inter-chiplet link behavior where the subcircuits connect, reaching fixed accuracy with less simulation time. SEQC’s stratification stage was inspired by CutQC, but unlike these works SEQC does not perform any classical “stitching” or simulation. It is interesting to note that the first version of SEQC was formulating the problem using constrained optimization as well, but the slowdown of the stratification step was unacceptable and impractical for large quantum circuits, and thus was abandoned in favor of the heuristics we currently employ in our compiler.
@@
Some works on qubit mapping share a few characteristics with SEQC. In a bottom-up approach to qubit mapping, Liu et al. (liu_2023_tackling) partition a circuit into blocks that span an identical number of qubits, perform permutation-aware synthesis on each block in parallel (solving for all possible qubit topologies), and finally use an augmented SABRE algorithm to map the blocks to the target device and route between blocks. SEQC takes a top-down approach instead, where SWAP permutations between chiplets are settled prior to any compilation tasks performed for each chiplet (block). As a result, SEQC only needs to solve for one qubit layout per block, rather than all permutations. Hungarian Qubit Assignment (HQA) (escofet2023hungarian) extends the Hungarian algorithm with the ability to adjust a qubit’s cost matrix according to behavior multiple time slices ahead. Pastor et al. (pastor2024circuit) use deep reinforcement learning to partition quantum circuits for execution across multiple chiplets. Both works aim to reduce inter-chiplet swaps, but assume that all inter-chiplet communications have the same cost, i.e., chiplets are fully connected, which is both impractical for large-scale systems and misses optimization opportunities. SEQC, in contrast, can work with arbitrary chiplet topologies. Ovide et al. (ovide2023mapping) leverage both quantum and classical methods for inter-chiplet communication to address the scalability challenges of monolithic processors. While it advocates for a two-level, hierarchical approach for quantum circuit mapping, it leaves the actual problem of qubit placement and routing unsolved. SEQC provides an algorithm and implementation, addressing both challenges.
Parallelizing SEQC’s recurrent elaboration step was inspired by traditional classical compilation techniques, which has a decades-long history and is now commonplace (Mickunas_1978_ParallelCompilation; Gross_1989_ParallelCompilation). Besides traditional compilation, though, we can also draw parallels between our work and prior research on data movement in the classical space. Data movement is widely understood to be the primary performance bottleneck in the classical architecture space. In the last two decades, work has proliferated on dataflow and spatial architectures, much of which focuses on mapping problems. Consider neural network architectures, for example, which seek to map their physical topologies as closely as possible to problem topologies. Eyeriss (Eyeriss_2016_Chen) approaches the dataflow mapping problem by proposing a row-stationary architecture, and then optimizing the mapping of problem to processing element such that elements which are spatially local in the problem are spatially local in hardware during execution. This is not unlike the target optimization for the stratification stage, and in fact is performed for many of the same reasons — like inter-chiplet swaps, data movements between more distant classical programming elements in a dataflow architecture are more expensive, and minimizing their frequency yields performance improvements.
@@ 8. Conclusion
Physical, technological, and economic considerations force future quantum systems to adopt a physically distributed architecture. Modular quantum processors already dominate the roadmaps of several major quantum companies. However, quantum compilers face a scalability challenge brought by the complexity and increasing scale of chiplet-based quantum devices, and their lack of chiplet awareness impedes their ability to produce high-quality results.
In this work, we first modify stock contemporary compilers to become aware of hardware modularity, thereby allowing them to correctly compile circuits for modular architectures with limited cross-chiplet gate support. We then propose SEQC, a Stratify-Elaborate Quantum Compiler for modular architectures. SEQC performs compilation in two stages, with the first stage (stratification, or chiplet splitting) performed only once for a given architecture, and the second stage (elaboration, or chiplet compilation) performed in parallel for each chiplet. Unlike contemporary quantum compilers that have to compile a quantum circuit anew before each execution to capitalize on the latest qubit calibration results, only the elaboration stage of SEQC needs to run before each execution. As its computational complexity is a function of the quantum chiplet size, not the entire quantum processor size, and chiplet sizes grow slowly, SEQC’s recurrent compilation time is largely unaffected by the relentless growth of qubit counts in future quantum processors.
SEQC implements several novel methods for qubit placement, qubit routing, and circuit optimization, with both hardware modularity and compilation parallelization in mind, achieving on average $9.3\%$ higher circuit fidelity (up to $49.99\%$), while attaining on average 3.27×3.27$\times$3.27 × faster compilation (up to 6.74×6.74$\times$6.74 ×) compared to a chiplet-unaware Qiskit baseline.

### v4

@@ 5. Experimental Results
Figures 11(b) and 12(b) show the recurring compilation time. Specifically, this refers to the elaboration time of SEQC and the full compilation time for Qiskit and PA-Qiskit. Overall, SEQC compilers have a large speedup over Qiskit-based compilers, owing to the use of parallelism and modular compilation strategies. Compared to Qiskit on the heavy-hex backend, SEQC-STA achieves $3.27\times$ speedup on average (up to $6.74\times$), and compared to PA-Qiskit on the heavy-hex backend, $4.69\times$ average speedup ($6.61\times$ max). Similarly on the grid backend, SEQC-STA achieves $1.34\times$ speedup on average (up to $3.37\times$) over Qiskit, and $2.52\times$ speedup on average (up to $11.93\times$) over PA-Qiskit.
(a) Compilation time.
(b) Relative recurring compilation time.
(c) Relative peak memory utilization.
Figure 11. Compilation performance metrics: compilation time, recurring compilation time, and peak memory utilization, 10-qubit heavy-hex backend (lower is better).
(a) Compilation time.
(b) Relative recurring compilation time.
(c) Relative peak memory utilization.
Figure 12. Compilation performance metrics: compilation time, recurring compilation time, and peak memory utilization, 120-qubit grid backend (lower is better).
Memory Utilization.
Figures 11(c) and 12(c) presents the relative peak memory utilization of the evaluated compilers. Compared to Qiskit, SEQC-SA exhibits $55.68\%$ higher memory usage on average ($117.71\%$ max) on the heavy-hex backend and $23.71\%$ higher memory usage on average ($38.24\%$ max) on the grid backend. This could likely be ameliorated by reducing the number of annealing trials, at the expense of circuit quality. SEQC-STA tends to use less memory compared to the Qiskit variants, by $4.22\%$ on average ($17.47\%$ max) on the heavy-hex backend and $10.69\%$ on average ($17.51\%$ max) on the grid backend. It is possible that the process of stratification and parallelized elaboration is responsible for the conservation of memory usage, presuming the memory requirements of quantum compilation scales super-linearly with the circuit size. This points to the scalability benefits of the SEQC framework in general.
@@ 6. Discussion
These results demonstrate a flaw with gate count and circuit depth as heuristics: their lack of noise-awareness. Circuit depth and gate count only consider the number of gates (in the critical path and in total) irrespective of each one’s unique underlying properties. It is expected that two-qubit gates will be significantly more noisy than single-qubit gates, and inter-module gates should be significantly more noisy than intra-module gates (smith_2022_chiplets). Failure to account for duration and error differences between gates or types of gates in quantum compilation heuristics will inevitably result in sub-optimal solutions when the discrepancies are extreme, e.g., in modular architectures. Conversely, the hierarchical structure of the SEQC framework with the stratification and elaboration stages imparts some degree of implicit noise-awareness, where minimizing the number of inter-module gates takes precedence over intra-module/total gates. As a result, SEQC can better optimize for the practically significant metrics of circuit fidelity and execution time.
Overall, we observe that SEQC yields circuits with more desirable characteristics when executing on modular quantum architectures, while at the same time it achieves faster compilation times. Also, as the number of qubits increases, the performance improvement of SEQC widens compared to the baseline, especially for the most important metrics of fidelity, compiled circuit execution time, and number of inter-module gates. These trends hold promise in allowing for faster and better compilation for future large-scale, hybrid modular quantum devices than we can achieve with today’s frameworks.
It is also important to note that SEQC is equally applicable to NISQ as well as future quantum systems. The stratification stage can decompose a large circuit into subcircuits that will be mapped to QPUs at the logical level, and then the elaboration stage can compile for each QPU together with an error correction method for that QPU. The two stages are distinct, and stratification operates more on the logical level, while the elaboration can be tasked with lowering from the logical to the physical qubits. Applicability to NISQ systems is still important, as even with technological breakthroughs, the transition to Fault Tolerant Quantum Computing (FTQC) will be gradual, not instantaneous. The hardware that exists now will continue improving and will be widely used for years, as there is practical scientific and industrial value in NISQ, even without the need to wait for fully error-corrected systems (piccinelli_chemistry; Kim2023preFTUtility; sciadv_utility_2025; Abanin2025ConstructiveInterference). Moreover, early fault-tolerant demonstrations will run on NISQ hardware, and will heavily rely on NISQ-era methods, as only few qubits can be protected. Thus, it remains critical to design architectures and compilers that serve NISQ hardware, both to extract near-term value, and to pave the engineering path toward full FTQC.
@@ 7. Related Work
The state-of-the-art in distributed quantum computing (DQC) (autocomm; 10228915; 10.1145/3730585; 10821229; 10313645; 10.1145/3579367) requires the generation and distribution of EPR pairs as the main entanglement resource over long-distance optical links, and the deployment of emitter-emitter, emitter-scatterer, or scatterer-scatterer photon-mediated entanglement schemes. However, modular architectures have less restrictive distance constraints, and they can often support other hardware technologies, such as microwave links (PhysRevApplied.21.054063). As a result, SEQC can model inter-module allocation topologically (as SWAP chains) (LaRacuente2025modelingshortrange), similar to compilation for monolithic (single-die) quantum processors. An analogy can be drawn to classical compilation between chiplet-based classical systems and distributed classical systems, where diverse compilation strategies are required due to very different hardware technologies and distances between computation units.
Compilation for modular transmon systems shares many similar constraints with compilation for quantum charge-coupled devices (QCCDs). Namely, these devices features modules, in the form of ion traps, and support movement between traps, via ion shuttling mechanisms (ovide2024sta). However, there are some key differences between the hardware technologies that make it difficult to compare SEQC with QCCD compilers on equal footing. Notably, qubits in superconducting devices needs to be topologically adjacent to each another to support two-qubit operations. However, QCCD devices are less restrictive, and support two-qubit operations so long as qubits are in the same traps / interaction region. As a result, QCCD compilers (10.1145/3526241.3530366; Kreppel2023quantumcircuit) can employ simpler routing strategies, or forgo them entirely, compared to modular transmon compilers.
@@
MECH (mech) entangles ancillary qubits and distributes them throughout a module to build fast network-on-chip routing channels (a ”communication highway”). MECH includes a compiler that requires, and optimizes for, the “highway” design. Unfortunately, the highway does not generalize to non-universal inter-module links. Future work to extend SEQC with MECH’s methodology may enable combining MECH’s techniques with SEQC’s allocation procedures to gain the benefits of each. Lin et al. (let_every_qubit) present a compilation technique that utilizes calibration data to determine potentially nonstandard 2Q gates that can perform standard operations along bases more compatible with their respective physical states and improve fidelity. Bandic et al.(bandic2024profiling) utilize the relationship between quantum circuit structure and mapping efficiency to optimize the architecture of monolithic and modular quantum processors, on the path to application-specific hardware optimization. Escofet et al. (escofet2024revisiting) develop a characterization technique for arbitrary quantum circuits, which provides theoretical upper/lower optimization bounds, enabling the optimality assessment of qubit mapping solutions. These techniques are orthogonal to SEQC, and could work synergistically with it.
@@
CutQC (cutqc) partitions large quantum circuits into smaller subcircuits that fit on a device, executes them one at a time, and then merges the individual results through classical post-processing. Piveteau and Sutter (circuit_knitting) similarly allow subcircuits to execute independently, and use quasiprobability simulation to simulate inter-module link behavior where the subcircuits connect, reaching fixed accuracy with less simulation time. SEQC’s stratification stage was inspired by CutQC, but unlike these works SEQC does not perform any classical “stitching” or simulation. We initially formulated qubit-to-subcircuit mapping as a constraint optimization problem, similar to CutQC, but existing numerical solvers were impractically slow for circuits larger than 20 qubits, corroborating the experience of Baker et al. (baker2020). Instead, we opted to use the approximate, but practical, solutions we currently employ in SEQC.
@@
Liu et al. (liu_2023_tackling) partition a circuit into blocks that span an identical number of qubits, perform permutation-aware synthesis on each block in parallel (solving for all possible qubit topologies), and finally use an augmented SABRE algorithm to map the blocks to the target device and route between blocks. SEQC takes a top-down approach instead, where SWAP permutations between modules are settled prior to any compilation tasks performed for each module (block). As a result, SEQC only needs to solve for one qubit layout per block, rather than all permutations.
Hungarian Qubit Assignment (HQA) (escofet2023hungarian) extends the Hungarian algorithm with the ability to adjust a qubit’s cost matrix according to behavior multiple time slices ahead. Pastor et al. (pastor2024circuit) use deep reinforcement learning to partition quantum circuits for execution across multiple modules. Both works aim to reduce inter-module swaps, but assume that all inter-module communications have the same cost, i.e., modules are fully connected, which is both impractical for large-scale systems and misses optimization opportunities. SEQC, in contrast, can work with arbitrary module topologies.
Ovide et al. (ovide2023mapping) leverage both quantum and classical methods for inter-module communication. While it advocates for a two-level approach to quantum circuit mapping, it leaves the actual problem of qubit placement and routing unsolved, which SEQC addresses with its algorithm.
@@ 8. Conclusion
[…]